OpenStudy (anonymous):

Pre-Alg. Chapter 1 Question! Mark plans to put concrete on a rectangular portion of his driveway. The portion is 8 feet long and 4 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Mark needs is $58.07. Which of the following is closest to the width of the portion of the driveway on which Mark plans to put concrete?

OpenStudy (anonymous):

(;_;)

OpenStudy (anonymous):

Hi! I exist!

OpenStudy (anonymous):

(u_u)

OpenStudy (anonymous):

Wow. I'm getting ignored so much that I could put "terms and conditions" out of business. @Vocaloid can you help me?

Vocaloid (vocaloid):

convert the height and length to yards first

Vocaloid (vocaloid):

still there?

OpenStudy (anonymous):

sorry my computer glitched up~ what do I convert to yards? 8ft long and 4in high?

OpenStudy (anonymous):

allo? ._.

OpenStudy (anonymous):

(~_~)

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

It would help to draw a picture. Use the [Draw] button, below.

OpenStudy (anonymous):

okey

Vocaloid (vocaloid):

yes, convert the dimensions to yards

Vocaloid (vocaloid):

convert 8 feet to yards convert 4 inches to yards

OpenStudy (anonymous):

4in = .11repeating yards. 8ft = 2.66667 yards now what?

OpenStudy (tkhunny):

 How do you calculate the volume of such a slab of concrete?

OpenStudy (anonymous):

volume? I dunno

OpenStudy (tkhunny):

I don't believe that. Given Length, Width, and Height, what is the volume of the Rectangular Prism?

OpenStudy (anonymous):

Wait, they gave the width?

OpenStudy (tkhunny):

No, they didn't. Just call it "Width". This is what the problem statement is asking. We must find it. Just write down the formula as if we know nothing, then fill in the pieces we know.

OpenStudy (anonymous):

Oh okay. So is it length x height = width?

OpenStudy (tkhunny):

Where did the volume go? Length * Width * Height = Volume Fill in the ones we know.

OpenStudy (anonymous):

Oh. So 2.6 * .11 * w = volume?

OpenStudy (tkhunny):

Try not to round so much. 8 ft = 2 2/3 yds = 8/3 yds 4 in = 4/36 yds = 1/9 yds We have (8/3)(1/9)(Width)=(Volume) Since we are given neither volume nor width, are we stuck? Reading the problem statement again, we see: "$98 per cubic yard. The total cost of the concrete Mark needs is $58.07." Can we get volume out of that information?

OpenStudy (anonymous):

Is the width less than a cubic yard? Because the total cost is less than the cost of the cubic yard

OpenStudy (tkhunny):

That's very good. We can use a proportion. If 1 yd^3 costs $98.00, how many yds^3 creates $58.07? Can you set up the proportion and solve for the necessary yds^3?

OpenStudy (anonymous):

I don't think I know how to do that.

OpenStudy (tkhunny):

BTW -- "Is the width less than a cubic yard?" Don't ever say that again. :-) "width" is a linear measurement. "cubic yard" is a volume measurement. They are not particularly related. They certainly can't be equal, greater then, or less than. It's like comparing 'x' to 'x^3'. You wouldn't add those, would you? Okay, "Like Terms" side-bar concluded.

OpenStudy (tkhunny):

Define C = Number of Cubic Yards of Concrete Then set up the proportion, \(\dfrac{1\;yd^{3}}{$98.00} = \dfrac{C\;yd^{3}}{$58.07}\) Can you solve for "C"?

OpenStudy (anonymous):

Not sure how

OpenStudy (tkhunny):

Multiply both sides by $58.07. This will do two things: 1) Get rid of everything but C on the RHS, and 2) Produce the result in the correct units on the LHS.

OpenStudy (anonymous):

How do you multiply fractions?

OpenStudy (tkhunny):

You should know this. You WILL be required to do it on a test. Right-Hand Side \($58.07\cdot\dfrac{C\;yd^{3}}{$58.07} = C\;yd^{3}\cdot\dfrac{$58.07}{$58.07} = C\;yd^{3}\cdot 1 = C\;yd^{3}\) Left-Hand Side \($58.07\cdot\dfrac{1\;yd^{3}}{$98.00}= 1\;yd^{3}\cdot\dfrac{$58.07}{$98.00}\) -- You finish that. It's just a division problem for your calculator.